1. If 5 times the value of a number is increased by 3, the result is 28. What is the number?
2. If 4 times the reciprocal of a number is 3 more than 5/2 times the reciprocal of that number, find the number.
3. John had $30,000 to invest. He invested part of this money in bonds paying 12% annual simple interest and the rest of the money in a savings account giving 4% annual interest. At the end of the year, he received $2,400 as extra income. How much money did John place in each investment?
4. The length of a football field is 180 feet more than its width. If the perimeter of the field is 1,060 feet, find the length of the field.
5. At a school choir concert, 256 students are standing in rows. If the number of students in each row is equal to the total number of rows, find the number of students in each row.
6. If 6 times a number is decreased by 8, the result is 40. What is the number?
7. Merry had $20,000 to invest. She invested part of this money in bonds paying 10% annual simple interest and the rest of the money in a savings account giving 5% annual interest. At the end of the year, she received $1,800 as extra income. How much money did Merry place in the savings account?
8. The length of a rectangular field is 50 feet less than its width. If the perimeter of the field is 840 feet, find the length of the field.
9. Alan participated in a car race in which he had to cover a distance of at least 50 kilometers. He had fuel in his car for a maximum distance of 53 kilometers. If the distance is given by , S(T)= 3T+ 47 where t is the time in hours, find the minimum and maximum number of hours for which Alan can drive his car.
10.In a 2-digit number, the number in the unit’s place is 4 more than the number in the ten’s place. If the digits of the number are reversed, the new number is 6 more than thrice the original number. Find the original number.

Answer by psbhowmick(878)

10. Let the unit's digit and ten's digit of the two digit number be X and Y respectively.

The number in the unit’s place is 4 more than the number in the ten’s place.
X = Y + 4 __________ (1)

If the digits of the number are reversed, the new number is 6 more than thrice the original number.
10X + Y = 3(10Y + X) + 6
10X + Y = 30Y + 3X + 6
7X - 29Y = 6 ____________ (2)

Substituting X = Y + 4 from equation (1) into (2)
7(Y + 4) - 29Y = 6
7Y + 28 - 29Y = 6
22Y = 22
Y = 1
X = Y + 4 = 5

So the two digit number is 15.
Contact me through email for solution to the other questions.